Six forces lying in a plane and forming angles of 60 relative to one another are applied to the centre of homogeneous sphere with mass M is equals to 6 kg. These forces are radially outward and consecutively 1N,2N,3N,4N,5N and6N. The acceleration of s p h e r e is
 

Grade:12

2 Answers

Hemanth
5 years ago
Hello Kkr,
here is the answer,
As all the forces are consecutive and are at angle of 60 degrees, They form as the sides from the center to the vertices of a regular hexagon.
If you do so , you’ll find that the vectors of magnitude (1,4), (2,5) and (3,6) are opposite (aligned at 180 degrees)
Finding their resultant, Three consecutive vectors of magnitude 3 and aligned at 60 degrees are formed.
Now at last finding their resultant , we get a vector of magnitude ‘6’ units
therefore, acceleration = Force /mass it implies
acceleration=6/6=1 m/s. That’s it
 
 
 
Khimraj
5 years ago
 
try to resolve force in opposite direction 
then resultant will have three force each of 3N at 0o, 60o, 120o
then resulatnt of it will have two force of 3N and 3\sqrt{3} at angle 90
then resultant will be \sqrt{(3\sqrt{3})^{2} + 3^{2}} = 6N
Then acceleration = 6/6 = 1m/s2
Hope it clears.................................